# xGen-MM (BLIP-3): A Family of Open Large Multimodal Models

### v2

@@ 5.2 Supervised Fine-tuning Data Recipe
The datasets used in the fine-tuning stage are from public datasets of different domains [29, 11, 49]. We sample data with various domain focuses including multi-modal conversation [29], image captioning [55, 56], visual question answering [57, 58, 59, 60], chart/document understanding [61, 62, 63, 64], science and math [65, 66]. In addition to the multi-modal image-text data, we also mix in pure text instruction following data  [67, 68] during visual instruction tuning. Ultimately, we collect a mixture of 1 million publically available instruction-tuning samples, on which we fine-tune our model for one epoch.
The multi-image instruction tuning stage starts with a model fine-tuned on single-image samples. We use a mixture of public multi-image/interleaved image-text instruction data [69, 70]. To prevent the model from deteriorating on single-image capabilities, we reuse a subset of single-image datasets used in the previous fine-tuning stage and mix them into the multi-image training data.
5.3 Post-training Data Recipe
Improving Truthfulness by Direct Preference Optimization.
We employ VLFeedback [71], a synthetically annotated multimodal preference dataset that uses off-the-shelf VLMs to generate responses to a diverse mix of multimodal instructions that are then scored by GPT4-V [2] along three axes – helpfulness, visual faithfulness, and ethics. The dataset contains 80k such instructions from which we construct preference data by marking as preferred (and dispreferred) the response with the highest (and lowest) average score across models and filtering out examples with low-scoring preferred responses. We thus generate 62.6k preference examples.
We perform 1 epoch of DPO on the combined preference dataset while updating a subset (2.5%) of LLM backbone weights using low-rank adaptation (LoRA [72]). Also, following recent work [50], we generate an additional set of responses that capture the model’s intrinsic hallucinations, by performing a second step of DPO per-iteration against the models’ output to a noised version of the input image and original query, which we treat as an additional dispreferred response.
Improving Harmlessness by Safety Fine-tuning.
Next, we perform 3 epochs of safety fine-tuning on the train split of the VLGuard [73] dataset, which contains 2k examples of unsafe images and instructions. VLGuard comprises two types of unsafe examples: (1) objectionable images paired with safe instructions and a desirable abstention response, and (2) safe images paired with two types of instruction-response pairs, one safe and another unsafe. The dataset consists of unsafe examples belonging to various subcategories including privacy-violating, risky/sensitive topics (such as politics, sex, and violence), deception, and discrimination. Following the original work, we randomly sample 5k additional examples from the instruction fine-tuning dataset to retain the model’s helpfulness without exaggerating its safety behavior. As before, we update a subset (2.5%) of LLM backbone weights using low-rank adaptation.
@@ 6 Experiments
6.1 Pre-training
Few-shot Evaluation.
After the pre-training stage, we evaluate our pre-trained model on classic captioning and VQA tasks, in comparison with previous models that support few-shot learning multi-modal evaluation. We present zero-shot and few-shot (4- and 8-shots) results, as shown in Table 1. Overall, our model achieves competitive multimodal in-context learning performance with comparable-sized LMMs  [^1] For few-shot evaluation, the zero-shot results are used mainly as a reference for their corresponding few-shot numbers, and they can be sensitive to prompts. As also mentioned in [9], they are mostly indicative of how well the pre-training distribution matches the associated evaluation task format. In pre-training evaluation, we mainly care for few-shot performance, which is robust to prompt templates.. For the OCR tasks (TextCaps and TextVQA) and VQA-v2, it significantly outperforms MM1-3B and even larger models such as Idefics-9B [13] and MM1-7B [9]. On all benchmarks, increasing the number of shots can improve the performance, demonstrating the model’s ability to adapt to in-context distributions.
Model  Shot Visual Question Answering Captioning
VQAv2 TextVQA OKVQA COCO NoCaps TextCaps
< 5B Model Comparisons
Flamingo-3B [22]  0 49.2 30.1 41.2 73.0
4 53.2 32.7 43.3 85.0
8 55.4 32.4 44.6 90.6
MM1-3B [9] 0 46.2 29.4 26.1 73.5 55.6 63.3
4 57.9 45.3 44.6 112.3 99.7 84.1
8 63.6 44.6 48.4 114.6 104.7 88.8
xGen-MM-base (4B) 0 43.1 34.0 28.0 67.2 82.6 69.5
4 66.3 54.2 48.9 107.6 100.8 89.9
8 66.9 55.3 50.1 109.8 104.6 94.0
Larger Models for Reference
Flamingo-9B [22] 8 58.0 33.6 50.0 99.0 - -
Idefics-9B [13] 8 56.4 27.5 47.7 97.0 86.8 63.2
MM1-7B [9] 8 63.6 46.3 51.4 116.3 106.6 88.2
Idefics2-8B [11] 8 70.3 57.9 54.6 116.0 - -
Table 1: Few-shot Pretraining Evaluation. Following [9], we randomly sample demonstrations from the training set as few-shot examples. We report CIDEr score for captioning and accuracy for VQA.
6.2 Supervised Fine-tuning
We evaluate our model on a comprehensive suite of multi-modal (image-text) benchmarks, assessing the model’s ability from multiple perspectives. Our evaluation covers general VQA benchmarks [74, 75, 76, 77, 78], visual perception [49], domain knowledge [79, 80], OCR ability [57, 81], and hallucination [82, 83]. For models fine-tuned on interleaved multi-image datasets, we also evaluate their performance on common multi-image benchmarks [69, 84, 85, 86].
Single-Image Evaluation.
In Table 2, we compare with models in comparable sizes ($<5$B), including both closed-source [9] and SoTA open-source models [10, 37]. We report individual benchmark scores along with two average scores: “Avg.(all)" is simply the average over all benchmarks, and “Avg.(perc.)" is the average score over benchmarks that focus on general VQA and visual perceptions. xGen-MM-instruct outperforms previous baselines on both general VQA and visual perception benchmarks. In addition, we find that xGen-MM-instruct-interleave, although further fine-tuned on multi-image data, maintains good performance on single-image benchmarks and has the highest overall scores.
Model (Size)
SEED
-IMG
SEED
v2
MMB
MM
Star
MME
CVB
-2D
CVB
-3D
RealW
QA
MMMU
Math
Vista
Sci
QA
@@ 6.2 Supervised Fine-tuning
Text
VQA
Avg.
Avg.
Closed-source models
GPT-4V 72.0 - 80.8 49.7 63.3 64.3 73.8 56.5 53.8 48.2 82.1 75.4 - - -
[…]
Open-source models
HPT-1.5-edge (4B) 72.3 - 74.6 45.8 - - - - 42.6 45.1 85.4 91.0 - - -
VILA-1.5-3B (3B) 67.9 - 63.4 - - - - - 33.3 - 69.0 85.9 - - -
VILA-1.5-3B (3B) 67.9 51.9 62.4 40.3 58.5 50.1 60.3 53.3 34.1 30.6 68.9 86.9 58.1 55.6 59.1
phi-3-vision (4B) - - 80.5 - - - - - - 44.5 90.8 85.8 70.9 - -
[…]
xGen-MM-inst.
-interleave (4B)
[…]
@@
Table 2: Evaluation on single-image benchmarks. phi-3-vision and VILA-1.5-3B are tested with our evaluation code [^3] https://github.com/open-compass/VLMEvalKit for a fair comparison. We also include the GPT-4V (gpt-4-1106-preview) performance (provided by the evaluation codebase) as a reference in the first row.
Multi-Image Evaluation.
In Table 3, we compare xGen-MM-instruct with xGen-MM-instruct-interleave on multi-image benchmarks. Although the former is fine-tuned from xGen-MM-base which can comprehend interleaved image-text data, it performs poorly on multi-image benchmarks. We suspect it is because solely fine-tuning on single-image data hurts such ability. With multi-image SFT, we see significantly improved scores. In addition, we also evaluate Xgen-MM-interleave on single-image benchmarks (See Table 2) and find that it maintains good performance on all benchmarks with the highest overall scores.
Model BLINK QBench-2 Mantis-eval
GPT-4V 51.1 73.4 62.7
VILA-1.5-3B (3B) 39.8 51.7 41.9
xGen-MM-inst. (4B) 46.6 52.4 42.4
xGen-MM-inst.-interleave (4B) 49.7 75.1 56.7
@@
Table 3: Evaluation on multi-image benchmarks. VILA-1.5-3B results are obtained using the same evaluation code as our models. We include the GPT-4V performance as a reference in the first row.
Figure 6: Example model outputs of xGen-MM-instruct-interleave. The model is capable of understanding interleaved image-text input and user queries about multiple images while maintaining the performance on single-image QAs.
6.3 Post-training
Table 4 summarizes the results of two post-training strategies for xGen-MM-instruct. We measure safety performance by ASR(attack success rate) on the VLGuard test split and hallucination performance using HallusionBench [82] (accuracy on image-context reasoning) and POPE [83] (average F1 score on binary entity presence questions). To ensure post-training doesn’t compromise helpfulness, we report performance on a few comprehension benchmarks as a control.
DPO enhances truthfulness by improving hallucination benchmarks (Row 2), while safety finetuning significantly reduces ASR (Row 3). Helpfulness is also improved slightly, as shown by control benchmarks. The final model, xGen-MM-dpo, includes both improvements.
Method Safety Hallucination Helpfulness (Control)
VLGuard ($\downarrow$) HalBench ($\uparrow$) POPE ($\uparrow$) SEED-IMG ($\uparrow$) MMB-dev ($\uparrow$) MME ($\uparrow$) MMStar ($\uparrow$)
xGen-MM-inst. (4B) 56.6 56.3 87.0 71.8 76.0 63.8 46.7
+ DPO 54.9 57.1 87.0 71.9 76.4 63.0 47.1
      + Safety FT 5.2 56.6 86.8 72.1 76.4 64.4 47.1
Table 4: Post-training results. We report results on safety and hallucination benchmarks after post-training, as well as on four helpfulness benchmarks as a control. Post-training improves harmlessness without compromising helpfulness.
@@ 7 Ablation Studies
7.1 Pre-training Ablation
Scaling Pre-training Data.
We perform an ablation study to explore the relation between the amount of pre-training data and the pre-train evaluation metrics, by varying the data scale from 2B multimodal tokens to 100B multimodal tokens. The data recipe we used here is a mixture of image caption datasets and multimodal interleaved data. As shown in Figure 7, we find that scaling up the number of multimodal tokens from 2B to 60B leads to substantial gain for image-text (COCO-Caps) and OCR (Text-Caps, TextVQA) tasks, and further increasing the data size to 100B has moderate additional benefit in terms of few-shot evaluation metrics.
@@ 7.1 Pre-training Ablation
(c) OK-VQA
Figure 7: Few-shot performance given different sizes of pretraining data.
Pre-training Data Recipe.
We discuss the impact of different data recipes for pre-training. Specifically, we perform ablation studies on top of a base data recipe: use Obelics [13] as the multimodal interleaved data source while keeping the caption datasets mixture the same. We also consider two other recipes (1) using MINT-1T [12] as interleaved data replacement, and (2) mixing additional pure text-only instruction-tuning data as a pre-train dataset. As shown in Table 5, we see a performance improvement using MINT-1T for image-text alignment (COCO-Caps) and OCR (Text-Caps, TextVQA), with a slight performance drop on OK-VQA, which is a knowledge-intensive task. We also find that adding text data can help attain the performance on OK-VQA that relies more on LLM capacity.
Data Text-VQA OK-VQA COCO-Caps Text-Caps
Obelics 41.1 / 41.9 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
MINT-1T 41.0 / 42.0 46.5 / 48.2 109.3 / 111.4 80.3 / 82.0
MINT-1T + text data 42.1 / 42.6 48.3 / 49.7 108.0 / 110.2 77.0 / 79.9
Table 5: Few-shot (4-shot / 8-shot) performance given different data recipes.
@@
We also explore if different visual backbones have an impact on the performance of vision-language tasks. We compare two types of visual encoders, DFN and SigLIP. Empirically, we find SigLIP provides better visual representations that boost performance on OCR tasks.
@@
Table 6: Few-shot (4-shot / 8-shot) performance given different visual backbones.
Number of Visual Tokens.
Another ablation is to study the impact of different numbers of visual tokens, i.e., input image tokens fed into the language model. We find that reducing the number of visual tokens from 128 to 64 can still attain similar performance, as shown in Table 7. This makes it possible for models to take in more visual images given a fixed context window.
Visual Token Text-VQA OK-VQA COCO-Caps Text-Caps
128 41.1 / 41.8 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
64 41.2 / 42.6 47.6 / 48.3 108.0 / 109.3 79.5 / 81.6
Table 7: Few-shot (4-shot / 8-shot) performance given the different number of visual tokens.
7.2 SFT Ablation
We conduct ablation studies at the instruction fine-tuning stage, focusing on several model design choices and data recipes. The SFT ablation studies are conducted on a simplified SFT data mixture, so the results in this section are not directly comparable to the main results in section  6.2.
Any-Resolution Vision Token Sampling.
@@ 7.2 SFT Ablation
Our any-resolution strategy differs from previous work [46] in that every group of image embeddings (of the same image patch) is downsampled with a perceiver resampler, which ensures that the number of vision tokens input to the LLM remains relatively small. In this section, we ablate the effectiveness of our any-resolution strategy by comparing it with a “fixed-resolution" baseline and other downsampling designs.
The “fixed-resolution" baseline resizes all images to the default input size of the vision encoder while keeping the original aspect ratios. We also tried another downsampling strategy with the perceiver resampler: Instead of doing downsampling for each patch independently, we consider a "fixed sampling" (denoted as anyres-fixed-sampling in Figure 8(a)). In the fixed sampling, we concatenate the image embeddings from all image patches and then input them as a single sequence to the perceiver resampler to obtain the fixed number of vision tokens for the whole image.
@@
Figure 8: SFT ablation studies. (a). Comparison of different vision token sampling strategies on OCR benchmarks. (b). Comparison between our model and its “instruction-aware" alternative. For each evaluation domain in Figure (b), we report the average score on multiple relevant benchmarks.
@@
Our evaluation of this design focuses on text-rich tasks (e.g., document understanding) that would benefit from high-resolution encoding with visual details. From Figure 8(a), we can see significant improvements with our resolution image encoding strategy even with downsampled vision tokens. The fixed sampling strategy, although it shows improvements over the base resolution baseline, is not as good as the patch-wise sampling. We suspect that this may be due to two reasons: (a) With fixed sampling, a vision token sequence that can have as long as over 3,000 embedding tokens will be compressed to 128 tokens, which may be too few to retain the information. (b) The perceiver resampler may not work well with a concatenation of different image embeddings.
Instruction-Aware Vision Token Sampling.
@@
InstructBLIP [7] proposes an instruction-aware Q-Former [1] for vision token sampling and shows that it can improve the model performance on some benchmarks. With the perceiver resampler as the VL connector, we can adopt a similar modification to make this process instruction-aware. To make our perceiver resampler “instruction-aware", we append the text instructions tokens to the query tokens of the perceiver resampler. Unlike Q-Former, there are only cross-attention layers inside the perceiver resampler, so the instruction (text tokens) would interact with both query tokens and image embeddings via cross-attention.
@@
From the comparison in Figure 8(b), we do not observe a significant difference between our model and its instruction-aware version on various benchmarks. It could be that our modification to the perceiver resampler can not be identical to the instruction-aware modification made to the Q-Former in  Dai et al. [7], and thus the effectiveness differs. Because of the little difference we observe in this ablation study, we keep the original perceiver resampler architecture in our model for simplicity. We leave the further exploration of the “instruction-aware" VL connector to future works.
Text-only SFT data
MMMU
MathVista
Science
QA
MME
MMStar
Conversation 39.1 37.1 84.8 64.9 46.1
Conversation + Math + Coding 40.9 38.9 81.4 64.8 45.3
Table 8: The impact of text-only SFT data. We compare two choices of text-only SFT data used for the image-text SFT data mixture.
Quality of the Text-only Instruction Data.
It is a common strategy to train or fine-tune a multi-modal LLM on both multi-modal and pure text data [8, 11]. It is mainly for maintaining the language ability of the fine-tuned model. In this experiment, we study how this pure text subset would affect the performance on multi-modal benchmarks. For the instruction tuning stage, we compare whether the diversity of the pure text data would affect the multi-modal performance. For example, how pure-text math data affects a model’s performance on multimodal math benchmarks. In our main experiments, the default collection of pure text instruction data covers diverse domains including conversation [67], math [68, 87], and code [88]. For this ablation study, we substitute these datasets with the same amount of samples that only cover general conversation.
In Table 8, we observe that adding math and coding data, although in pure-text format, can help improve a model on relevant benchmarks like MathVista [80], while has less effects on general VQA benchmarks.

### v3

@@ 5.2 Supervised Fine-tuning Data Recipe
The datasets used in the fine-tuning stage are from public datasets of different domains [48, 38, 71, 75, 12, 69, 61, 68, 31, 59, 58, 34, 35, 33, 56, 84]. In addition to the multi-modal image-text data, we also mix in pure text instruction-following data  [45, 62] during visual instruction tuning. Ultimately, we collect a mixture of 3 million publicly available instruction-tuning samples, on which we fine-tune our model for one epoch.
The multi-image instruction tuning stage starts with a model fine-tuned on single-image data. We use a mixture of public multi-image / interleaved image-text instruction data [32, 54]. To prevent the model from deteriorating on single-image capabilities, we reuse a subset of single-image datasets used in the previous fine-tuning stage and mix them into the multi-image training data.
@@ 6 Experiments
Model Accessibility
SEED-IMG
MMB(dev)
MMStar
MME(norm)
RWQA
MMVet
MMMU(val)
MathVista
TextVQA
OCRBench
@@
HalBench
Average
open-
weight
open-
data
open-
code
data
contribution
GPT-4V [63] $\times$ $\times$ $\times$ $\times$ 72.0 80.8 49.7 63.3 61.4 49.0 53.8 48.2 78.0 67.8 81.8 39.3 62.1
3B-5B Models
MM1.5-3B [82] $\times$ $\times$ $\times$ $\times$ 72.4 - - 64.2 56.9 41.0 37.1 44.4 76.5 65.7 88.1 - -
Phi-3-vision [1] $\times$ $\times$ $\times$ 71.0 73.6 47.9 55.3 58.8 43.2 46.1 45.1 73.3 63.7 83.5 39.0 58.4
Phi-3.5-vision [1] $\times$ $\times$ $\times$ 69.6 81.9 47.3 65.6 53.5 43.2 43.0 43.9 72.0 59.9 86.1 39.6 58.8
MiniCPM-V-2.0 [76] $\times$ $\times$ $\times$ 67.1 69.1 39.1 64.6 55.8 41.0 38.2 39.8 73.2 59.6 86.3 36.1 55.8
VILA-1.5-3B [47] $\times$ 67.9 62.4 40.3 58.5 53.3 38.5 34.1 30.6 58.1 43.7 86.9 31.2 50.5
BLIP-3-4B (ours) 71.9 74.7 45.3 62.5 61.0 41.2 42.3 44.5 76.5 66.0 87.9 36.5 59.2
10B-15B Models
Pixtral-12B [2] $\times$ $\times$ $\times$ 71.5 77.9 54.5 68.6 65.4 58.5 51.1 56.3 75.7 72.2 84.2 47.0 64.9
Llama-3.2-11B-V-inst. [55] $\times$ $\times$ $\times$ 72.7 68.0 49.8 65.0 63.3 57.6 48.0 47.7 54.1 75.3 88.1 40.3 60.8
VILA-1.5-13B [47] $\times$ 68.0 74.4 44.2 61.4 53.3 45.0 41.1 42.3 61.1 45.7 85.0 39.3 55.1
Cambrian-13B [71] 73.2 73.2 47.1 67.0 58.6 48.9 41.6 47.7 72.8 61.0 86.8 39.4 59.8
BLIP-3-14B (ours) 73.4 77.5 48.7 68.1 64.4 44.5 47.0 45.4 79.0 71.1 86.0 38.7 62.0
@@
Table 1: Evaluation on single-image benchmarks. For benchmarks where a open-source model doesn’t report its official score, we use the score reported in a third-party leaderboard [^2] https://huggingface.co/spaces/opencompass/open_vlm_leaderboardor run it ourselves using the evaluation codebase [18] for a fair comparison. We also include the GPT-4V (gpt-4-1106-preview) performance (provided by the evaluation codebase) as a reference in the first row.
We evaluate our models (4B and 14B) on a comprehensive suite of multimodal benchmarks, assessing the model’s ability from multiple perspectives. Our evaluation covers general VQA benchmarks [39, 52, 20, 13, 3], domain knowledge [80, 57], OCR ability [69, 53], and hallucination [27, 44]. For models fine-tuned on interleaved multi-image datasets, we also evaluate their performance on common multi-image benchmarks [32, 74, 72, 21].
Single-Image Evaluation. In Table 1, we compare models of comparable sizes, including both closed-source [63, 82] and open-source models [47, 1, 76, 2, 55, 71]. We report individual benchmark scores as well as the overall average score across all benchmarks, following the standard practice.
Model BLINK QBench MuirBench Mantis-Eval
GPT-4V [63] 51.1 73.4 - 62.7
3B-5B Models
VILA-1.5-3B [47] 39.8 51.7 29.2 41.9
BLIP-3-4B-SI 42.9 51.9 35.0 49.3
BLIP-3-4B-MI 47.1 69.6 38.0 56.2
10B - 15B Models
VILA-1.5-13B [47] 48.1 61.2 37.7 49.3
BLIP-3-14B-SI 47.1 55.4 49.1 55.3
BLIP-3-14B-MI 53.9 73.4 56.2 61.3
@@
Table 2: Evaluation on multi-image benchmarks. VILA-1.5 models are evaluated using the same evaluation code as our models.
Multi-Image Evaluation. In Table 2, we compare BLIP-3 single-image (SI) model with BLIP-3 interleaved multi-image (MI) model on multi-image benchmarks. Although the former is fine-tuned from a pre-trained model that can comprehend interleaved image-text data, it performs poorly on multi-image benchmarks, possibly because the single-image SFT causes degradation in such ability. With interleaved multi-image SFT, we see significant improvements.
@@ 7 Ablation Studies
7.1 Ablation on Stage-1 Pre-training
Few-shot Pre-training Evaluation. To analyze the impact of pre-training configurations without the impact of fine-tuning data, we conduct ablation studies on pre-trained models with few-shot evaluation. Following [60], we randomly sample a few-shot subset from the training set and evaluate model performance. We report CIDEr scores for captioning tasks and accuracy for VQA to provide a clear comparison of few-shot performance across different pre-training configurations.
Scaling Pre-training Data. For Stage-1 pre-training, we perform an ablation study to explore the relation between the amount of pre-training data and the pre-train few-shot evaluation metrics, by varying the data scale from 2B multimodal tokens to 100B multimodal tokens. The data recipe we used here is a mixture of image caption datasets and multimodal interleaved data as detailed in 3. As shown in Figure 5, we find that scaling up the number of multimodal tokens from 2B to 60B leads to substantial gain for image-text (COCO-Caps) and OCR (Text-Caps) tasks, and further increasing the data size to 100B has moderate additional benefit in terms of few-shot evaluation metrics.
@@ 7.1 Ablation on Stage-1 Pre-training
Figure 5: Few-shot performance given different scales of pre-training data in Stage-1 pre-training.
@@
We also explore if different visual backbones have an impact on the performance of vision-language tasks. We compare two types of visual encoders, DFN [19] and SigLIP [81]. As shown in 3, we find SigLIP provides better visual representations that boost performance on OCR tasks, and we adopt SigLIP in the final model architecture as the ViT backbone. To ensure computational efficiency and a fair comparison, all models in this ablation study are pre-trained with 10B tokens.
@@
Table 3: Few-shot (4-shot / 8-shot) performances for different visual backbones at Stage-1 pre-training.
7.2 Ablation on Stage-2 Pre-training
In this section, we investigate the impact of Stage-2 pre-training. Model performance is evaluated by fine-tuning on a smaller subset ($\sim$ 1M samples) of our instruction tuning dataset for better computational efficiency. For evaluation, we focus on two representative LMM capabilities: general VQA and OCR, each of which is an averaged score over multiple relevant benchmarks.
Impact of Adding Stage-2 Pre-training and with Different Resolutions. As shown in Table 4, we use our 100B token Stage-1 pre-trained model as the baseline. We then add Stage-2 pre-training on top of it; we use half the data samples for a full Stage-2 pre-training run to maintain computational efficiency. Furthermore, we examine how varying Stage-2 pre-training resolutions (i.e., different numbers of image patches) affect performance. To ensure a fair comparison, all models in Table 4 are fine-tuned with a maximum of 13 image patches and follow the same fine-tuning protocol. From the results, we observe that incorporating Stage-2 pre-training leads to a significant improvement in OCR, while general VQA performance remains stable.
Stage-2 Stage-2 Maximum Resolutions OCR GeneralVQA
$\times$ N/A 64.6 63.5
(768x768), (384x1152), (1152x384) 65.3 63.2
(1152x1536), (1536x1152) 67.2 63.3
Table 4: Stage-2 pre-training and its resolution impact. We ablate two resolution grid settings for Stage-2 pre-training: one with up to 4 image patches, at most arranged as (2×2, 1×3, or 3×1), and another with up to 12 image patches, at most arranged as (3×4 or 4×3), with each patch sized at 384×384 pixels.
Figure 6: Example model outputs of BLIP-3-4B-MI (interleaved multi-image model). The model is capable of understanding interleaved image-text input and user queries about multiple images while maintaining the performance on single-image QAs.
7.3 Ablation on Instruction Tuning
The ablation study in this subsection focuses on several model design choices. Evaluations are conducted on single-image fine-tuned checkpoints. Similar to section 7.2, models in this ablation are fine-tuned on a smaller SFT subset for better computational efficiency, so the results are not directly comparable to our main results.
Perceiver Resampler vs. MLP. We experiment with two designs of vision token sampler with similar “vision compression ratios”. Both models are trained from Stage-1 pre-training. For perceiver resample, we use 128 query tokens per image patch to achieve a compression ratio of $r_{c}=729/128=5.7$ With MLP projector, to get a similar compression ratio, we append a 2x2 max pooling layer to it to have $r_{c}=4$. In Table 5, we compare the two models on benchmarks focusing on various domains and report the average score on each domain. Perceiver resampler, with a slightly higher compression ratio, has a lower score on OCR, but performs better in other domains, and has a better overall score. We therefore choose to use perceiver resampler for BLIP-3, because the higher compression ratio is crucial for efficiency when handling interleaved multi-image input.
Vision Token Sampler GeneralVQA OCR Sci.& Math Hallucination
Perceiver resampler 63.1 70.4 43.4 62.2
MLP + 2x2 pooling 61.7 72.2 41.8 62.7
Table 5: Ablation study on vision token sampler. We experiment with two sampling designs with similar vision compression ratio.
@@ 7.3 Ablation on Instruction Tuning
Any-Resolution Vision Token Sampling. Our any-resolution strategy differs from previous work [50] in that every group of image embeddings (of the same image patch) is downsampled with a perceiver resampler to ensure efficiency. In this section, we ablate the effectiveness of our any-resolution strategy by comparing it with other designs. Models in this study are fine-tuned from the Stage-1 pre-trained base resolution models.
@@
Figure 7: SFT ablation studies. (a). Comparison of different vision token sampling strategies on OCR benchmarks. (b). Comparison between our model and its “instruction-aware” alternative. For each evaluation domain in Figure (b), we report the average score on multiple relevant benchmarks.
The “fixed-resolution” baseline resizes all images to the default input size of the vision encoder while keeping the original aspect ratios. Another downsampling strategy with the perceiver resampler is “fixed sampling” ( anyres-fixed-sampling), for which we concatenate the image embeddings from all image patches and then input them as a single sequence to the perceiver resampler to obtain the fixed number of vision tokens for the whole image.
@@
Our evaluation of this design focuses on text-rich OCR benchmarks. From Figure 7(a), we see significant improvements with the any-res image encoding strategy even with downsampled vision tokens. The fixed sampling strategy, although it shows improvements over the base resolution baseline, is not as good as the patch-wise sampling. We suspect two reasons for this: (a) With fixed sampling, the vision token compression ratio can be too high to retain text-rich visual information. (b) The perceiver resampler may not work well with a concatenation of different image embeddings.
@@
Instruction-Aware Vision Token Sampling. InstructBLIP [15] proposes an instruction-aware Q-Former [42] for vision token sampling and proves its effectiveness on certain benchmarks. With the perceiver resampler, we experiment with a similar strategy by appending text instruction tokens to the query tokens of the perceiver resampler. This enables the instruction (text tokens) to interact with both query tokens and image embeddings via cross-attention.
@@
From the comparison in Figure 7(b), we do not observe a significant difference between our model and its instruction-aware version on common task domains. Because of the little difference we observe in this ablation study, we keep the original perceiver resampler architecture in our model for simplicity. We leave the further exploration of the “instruction-aware” VL connector to future works.